# On the convergence of noisy Bayesian Optimization with Expected Improvement

### v1

@@ 4.2 Improved error bound
(105) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w)-\tau(z_{t_{k}})>0.$
From (91), (92) and (105), we again have (97). Therefore, (97) is valid for Case 2.1.1..
Case 2.1.2. Next, consider $z_{t_{k}}\leq-C_{3}$. By (90), $\frac{b_{t_{k}}}{b^{*}_{t_{k}}}\leq\frac{w}{C_{3}}<1$. Using the monotonicity of $\tau(\cdot)$, (91) implies
@@
That is, we have (97) for Case 2.1.2. Combining Case 2.1.1 and Case 2.1.2, we have $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})>EI(a_{t_{k}},b_{t_{k}})$ if $a_{t_{k}}<0$ (Case 2.1.). Figure 3 demonstrates an example contour of $\bar{\tau}(\rho;z,w)-\tau(z)$ for given $w$ and $C_{3}$.
Case 2.2. Next, we consider $a_{t_{k}}\geq 0$. By definition (85),
@@
Thus, $C_{3}\rho_{t_{k}}-w\leq 0$. Define function $\tilde{\tau}:\mathbb{R}^{2}\to\mathbb{R}$ as
(109) $\tilde{\tau}(\rho,z;w)=\frac{1}{\rho}\tau\left(C_{1}z\rho+C_{3}\rho-w\right),$
for $\rho\in(0,\frac{w}{C_{3}}),z\geq 0$. From (107), $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})\geq b_{t_{k}}\tilde{\tau}(\rho_{t_{k}},z_{t_{k}};w)$. We now compare $\tilde{\tau}(\rho_{t_{k}},z_{t_{k}};w)$ and $\tau(z_{t_{k}})$. For $\forall\rho_{t_{k}}\in(0,\frac{w}{C_{3}})$, (80) and Lemma 3.2 lead to the derivative
(110) $\frac{\partial\tilde{\tau}}{\partial z}=C_{1}\Phi\left(C_{1}z\rho+C_{3}\rho-w\right)>C_{1}\Phi(-w)>\Phi(z)=\frac{d\tau}{dz}.$
Therefore, both $\tilde{\tau}(\cdot,\cdot)$ and $\tau(\cdot)$ are monotonically increasing with $z>0$ and $\tilde{\tau}$ always has a larger positive derivative. For any $\rho_{t_{k}}$, this leads to
(111) $\tilde{\tau}(\rho_{t_{k}},z_{t_{k}};w)-\tau(z_{t_{k}})\geq\tilde{\tau}(\rho_{t_{k}},0;w)-\tau(0)=\frac{1}{\rho_{t_{k}}}\tau\left(C_{3}\rho_{t_{k}}-w\right)-\tau(0).$
@@
[…]
Given (80),
(114) $\bar{\tau}(\rho_{t_{k}};0,w)-\tau(0)>C_{3}\Phi(-w)-\phi(0)>0.$
@@
To summarize, under the Scenario B Case 2, for all cases given (90),
$EI(a_{t_{k}}^{*},b_{t_{k}}^{*})>EI(a_{t_{k}},b_{t_{k}})$. Combined with (89), under Scenario B, $EI(a_{t_{k}}^{*},b_{t_{k}}^{*})>EI(a_{t_{k}},b_{t_{k}})$ with probability greater than $1-\delta_{\sigma}$. Combined with (87), with probability greater than $1-\frac{2\delta}{9}-\delta_{\sigma}$, $EI_{t_{k}}({\bm{x}}^{*})>EI_{t_{k}}({\bm{x}}_{t_{k}+1})$. However, this is a contradiction of (77), a sure event. Therefore, we have
@@
That is, $\delta_{\sigma}\geq 1-\frac{2\delta}{9}$. From (84),
@@
Notice again that (117) and (76) both stand with probability greater than $1-\delta$, as $\delta$ is derived from (72), Lemma 3.15 and Lemma 3.12. Under Scenario B, by (117) and (76), with probability greater than $1-\delta$,
@@
From (80), $C_{1}>2$ and $C_{2}>1$. Combining Scenario A (82) and Scenario B (118), we obtain (70).
If there is no noise, i.e., $\sigma=0$, we have $f({\bm{x}}_{t})=y_{t}\geq y_{t}^{+}$. Therefore, we can choose $k=[\frac{t}{2}]$ and there exists $k\leq t_{k}\leq 2k$ so that $y_{t_{k}}^{+}-y_{t_{k}+1}^{+}<\frac{2M}{k}$ with probability $1$. Following the same proof procedure as the noisy scenario, we obtain (71).
4.3 Convergence rate and comparative analysis
The rate of decrease for the noiseless error bound of Theorem 4.7 is given below in the setup of [14]. The readers are referred to [14] for the underlying assumptions on the kernels and the choice of parameters $\nu>0$ and $\alpha\geq 0$.
@@ 4.3 Convergence rate and comparative analysis
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{3c_{\alpha}}{\delta})$, where $c_{\alpha}=\frac{1+2\pi}{\pi}$, $C_{1}=\frac{1}{\Phi(-\sqrt{\beta})}$ and $C_{2}=\frac{\phi(0)}{\Phi(-\sqrt{\beta})}+\sqrt{\beta}$. Suppose the kernels $k(\cdot,\cdot)$ satisfy assumptions 1-4 in [14], then there exists constant $C^{\prime}$ so that
(119) $\mathbb{P}\left\{r_{t}\leq 2C_{1}M\frac{3}{t-3}+(C_{1}\sqrt{\beta}+C_{2})C^{\prime}(\frac{t-3}{3})^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}\left(\frac{t}{3}\right)\right\}\geq 1-\delta.$
Hence, the convergence rate of GP-EI is $\mathcal{O}(t^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(t))$.
@@
By Lemma 7 in [14], there exists a constant $C^{\prime}>0$ so that for $\forall k\in\mathbb{N}$ and ${\bm{x}}_{i},i=1,\dots,t$, $\sigma_{i}({\bm{x}}_{i+1})\geq C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)$ holds at most $k$ times, where $\eta=\alpha$ if $\nu\leq 1$ and $\eta=0$ if $\nu>1$. Both $\nu>0$ and $\alpha\geq 0$ are parameters defining the properties of the kernels in assumptions 1-4 in [14].
Following the proof of Theorem 4.7, we can let $k=\left[\frac{t}{3}\right]$. Then, there exists $k\leq t_{k}\leq 3k$ so that $y_{t_{k}}^{+}-y_{t_{k}+1}<\frac{2M}{k}$ and $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})<C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)$. Applying Theorem 4.7, with probability $1-\delta$,
@@
Theorem 4.9 illustrates that the convergence rate in the noiseless case under the GP prior assumption is similar to that under the RKHS assumption in [14]. The analysis can be applied to Theorem 4.2 in the noiseless case as well, with a similar convergence rate.
The rate of decrease for the error bound with noise is given below.
@@
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{9c_{\alpha}}{\delta})$, $w=\sqrt{2\log(\frac{9}{2\delta})}$, and $c_{t}^{\sigma}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Let $C_{1}=\frac{1}{\Phi(-w)}$ and $C_{2}=\frac{\phi(0)}{\Phi(-w)}+\sqrt{\beta}$. Then, for $t\geq\frac{4\log(\frac{3}{\delta})}{\log(2)}+4$, the error bound of the EI algorithm reduces at the rate
@@
Next, we prove the error bound in Theorem 4.7 is smaller than that of Theorem 4.2.
@@
We prove the noisy case as the noiseless case is similar. The corresponding $\beta_{2}$ in Theorem 4.7 is $\beta_{2}=2\log(\frac{9c_{\alpha}}{\delta})$, and the $\beta_{1}$ in Theorem 4.2 is $\beta_{1}=2\log(\frac{6}{\delta})$. Clearly, the probabilities for (60) and (70) are the same. To better compare the two theorems, we can rewrite both  (60) and (70) as
@@
(124) $C_{4}^{1}=\frac{\tau(\sqrt{\beta_{1}})}{\tau(-\sqrt{\beta_{1}})},\ C_{5}^{1}=\frac{\tau(\sqrt{\beta_{1}})}{\tau(-\sqrt{\beta_{1}})}(\sqrt{\beta_{1}}+\phi(0)),$
where the superscript $1$ indicates Theorem 4.2 and $2$ refers to Theorem 4.7. Let $w_{2}=2\log(\frac{9}{2\delta})$. From (70),
(125) $C_{4}^{2}=\frac{1}{\Phi(-\sqrt{w_{2}})},\ C_{5}^{2}=\frac{\sqrt{\beta_{2}}}{\Phi(-\sqrt{w_{2}})}+\frac{\phi(0)}{\Phi(-\sqrt{w_{2}})}+\sqrt{\beta_{2}}.$
Using $\beta_{1}>w_{2}$, we have
(126) $C_{4}^{1}=\frac{\tau(\sqrt{\beta_{1}})}{\tau(-\sqrt{\beta_{1}})}>\frac{\tau(\sqrt{\beta_{1}})}{\Phi(-\sqrt{\beta_{1}})}>\frac{\sqrt{\beta_{1}}}{\Phi(-\sqrt{w_{2}})}>C_{4}^{2},$
where the first inequality uses Lemma 3.4, the second 3.6, and the last $\beta_{1}>1$. Notice that $w_{2}>2\log(\frac{9}{2})>3$. Then, using $\delta\in(0,1)$, $w_{2}<\beta_{2}$, and inequalities from (126), we have
(127) $C_{5}^{2}=$ $\frac{1}{\Phi(-\sqrt{w_{2}})}(\sqrt{\beta_{2}}+\phi(0)+\sqrt{\beta_{2}}\Phi(-\sqrt{w_{2}}))<\frac{1}{\Phi(-\sqrt{\beta_{1}})}(1.042\sqrt{\beta_{2}}+\phi(0)).$
It is easy to verify that $1.042\sqrt{\beta_{2}}+\phi(0)<\beta_{1}+\phi(0)\sqrt{\beta_{1}}$ by taking the square on both sides. Hence, by Lemma 3.4 and 3.6
(128) $C_{5}^{2}<$ $\frac{1}{\Phi(-\sqrt{\beta_{1}})}(\beta_{1}+\phi(0)\sqrt{\beta_{1}})<\frac{\tau(\sqrt{\beta_{1}})(\sqrt{\beta_{1}}+\phi(0))}{\tau(-\sqrt{\beta_{1}})}=C_{5}^{1}.$
@@
An example of Proposition 4.14 is $\delta_{1}=0.1$. Then, $\beta_{1}=8.19$, $C_{4}^{1}=4632$ and $C_{5}^{1}=15103$. Meanwhile, $\beta_{2}=9.17$, $C_{1}=345$, and $C_{2}=141$. Thus, $C_{4}^{2}=345$, and $C_{5}^{2}=1187$. An illustration of Proposition 4.14 is shown in Figure 5.
Remark 4.17.
Another improvement of an upper bound in the form of Theorem 4.7 compared to Theorem 4.2 is that the cumulative regret can be computed with a sum of $\sigma_{t}({\bm{x}}^{*})$. Then, following the proof of Theorem 4.7, the cumulative regret bounds may be established for GP-EI for the first time. We leave this topic for future investigation.
4.4 Improved error bound for RKHS objective
In this section, we demonstrate how the analysis and techniques from Section 4.1 and 4.3 can be applied to RKHS objective to improve existing best-known error bounds and convergence rates from [14] in the noiseless case. Formally, Assumptions 2.52.6 are valid in this section. Additionally, we maintain a small value of $\sigma$, e.g., $10^{-8}$ in computing the posterior mean and variance in (2), even though there is no observation noise. The following lemma can be found in [29].

### v2

@@ 4.1 Asymptotic convergence in the noisy case
(97) $\bar{\tau}(\rho_{t_{k}};z_{t_{k}},w,C_{3})-\tau(z_{t_{k}})>0.$
From (83), (84) and (97), we again have (89). Therefore, (89) holds for Case 2.1.1.
Case 2.1.2 Next, consider $z_{t_{k}}\leq-C_{3}$. By (82), $\frac{b_{t_{k}}}{b^{*}_{t_{k}}}\leq\frac{w}{C_{3}}<1$. Using the monotonicity of $\tau(\cdot)$, (83) implies
@@
That is, we have (89) for Case 2.1.2. Combining Case 2.1.1 and Case 2.1.2, we have $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})>EI(a_{t_{k}},b_{t_{k}})$ if $a_{t_{k}}<0$ (Case 2.1). Figure 3 demonstrates an example contour of $\bar{\tau}(\rho;z,w)-\tau(z)$ for some fixed $w$ and $C_{3}$.
Case 2.2 Next, we consider $a_{t_{k}}\geq 0$. By definition (77), we can write
@@
Thus, $C_{3}\rho_{t_{k}}-w\leq 0$. Define function $\tilde{\tau}(\rho,z):\mathbb{R}^{2}\to\mathbb{R}$ with parameter $w,C_{1},C_{3}$ as
(101) $\tilde{\tau}(\rho,z;w,C_{1},C_{3})=\frac{1}{\rho}\tau\left(C_{1}z\rho+C_{3}\rho-w\right),$
for $\rho\in(0,\frac{w}{C_{3}}),z\geq 0$. From (99), $EI(a^{*}_{t_{k}},b^{*}_{t_{k}})\geq b_{t_{k}}\tilde{\tau}(\rho_{t_{k}},z_{t_{k}})$. We now compare $\tilde{\tau}(\rho_{t_{k}},z_{t_{k}};w,C_{1},C_{3})$ and $\tau(z_{t_{k}})$. For $\forall\rho_{t_{k}}\in(0,\frac{w}{C_{3}})$, (74) and Lemma 3.2 lead to the derivative
(102) $\frac{\partial\tilde{\tau}}{\partial z}=C_{1}\Phi\left(C_{1}z\rho+C_{3}\rho-w\right)>C_{1}\Phi(-w)>\Phi(z)=\frac{d\tau}{dz}>0.$
Therefore, both $\tilde{\tau}(\cdot,\cdot)$ and $\tau(\cdot)$ are monotonically increasing with $z\geq 0$ and $\tilde{\tau}$ always has a larger positive derivative. For any $\rho_{t_{k}}$, this leads to
(103) $\tilde{\tau}(\rho_{t_{k}},z_{t_{k}};w,C_{1},C_{3})-\tau(z_{t_{k}})\geq\tilde{\tau}(\rho_{t_{k}},0;w,C_{1},C_{3})-\tau(0)=\frac{1}{\rho_{t_{k}}}\tau\left(C_{3}\rho_{t_{k}}-w\right)-\tau(0).$
@@
(105) $\bar{\tau}(\rho_{t_{k}};0,w,C_{3})-\tau(0)>$ $\frac{1}{\bar{\rho}}\tau(C_{3}\bar{\rho}-w)-\phi(0)=\frac{1}{\bar{\rho}}((C_{3}\bar{\rho}-w)\Phi(C_{3}\bar{\rho}-w)+$
$\phi(C_{3}\bar{\rho}-w))$ $-\phi(0)=C_{3}\Phi(C_{3}\bar{\rho}-w)-\phi(0)>C_{3}\Phi(-w)-\phi(0).$
Then, (74) implies that
(106) $\bar{\tau}(\rho_{t_{k}};0,w,C_{3})-\tau(0)>C_{3}\Phi(-w)-\phi(0)=0.$
@@
To summarize Scenario B Case 2, if (82), then $EI(a_{t_{k}}^{*},b_{t_{k}}^{*})>EI(a_{t_{k}},b_{t_{k}})$. Combined with (81), under Scenario B, $EI(a_{t_{k}}^{*},b_{t_{k}}^{*})>EI(a_{t_{k}},b_{t_{k}})$ holds with probability greater than $1-\delta_{\sigma}$. Combined with (79), we have with probability greater than $1-\frac{2\delta}{9}-\delta_{\sigma}$ that $EI_{t_{k}}({\bm{x}}^{*})>EI_{t_{k}}({\bm{x}}_{t_{k}+1})$. However, this is a contradiction of (69), which is a sure event. Therefore, we have
@@
That is, $\delta_{\sigma}\geq 1-\frac{2\delta}{9}$. Then, (76) implies that
@@
We note again that (109) and (68) hold simultaneously with probability greater than $1-\delta$, as $\delta$ is derived from (64), Lemma 3.15, and Lemma 3.12. Under Scenario B, by (109) and (68), we obtain with probability greater than $1-\delta$ that
@@
From (74), we know $C_{1}>2$ and $C_{2}>1$. Combining Scenario A (73) and Scenario B (110), one can obtain (62).
If there is no noise, i.e., $\sigma=0$, we have $f({\bm{x}}_{t})=y_{t}\geq y_{t}^{+}$. Therefore, we can choose $k=[\frac{t}{2}]$ and there exists $k\leq t_{k}\leq 2k$ so that $y_{t_{k}}^{+}-y_{t_{k}+1}^{+}<\frac{2M}{k}$ with probability $1$. Following the same proof procedure above in the noisy case, we obtain (63).
4.2 Convergence rates and comparative analysis
The rate of decrease for the noiseless error bound of Theorem 4.7 is given below in the setup of [17]. Recall that $\nu$ is the Matérn kernel parameter. Define $\eta=\begin{cases}\alpha,\ \nu\leq 1\\ 0,\ \nu>1,\end{cases}$ where $\alpha=\frac{1}{2}$ if $\nu\in\mathbb{N}$, and $\alpha=0$ otherwise. The readers are referred to [17] for the underlying assumptions on the kernels and the choice of parameters $\nu>0$ and $\alpha\geq 0$.
@@ 4.2 Convergence rates and comparative analysis
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{3c_{\alpha}}{\delta})$, where $c_{\alpha}=\frac{1+2\pi}{\pi}$, $C_{1}=\frac{1}{\Phi(-\sqrt{\beta})}$ and $C_{2}=\frac{\phi(0)}{\Phi(-\sqrt{\beta})}+\sqrt{\beta}$. Suppose the kernels $k(\cdot,\cdot)$ satisfy assumptions 1-4 in [17]. Then, there exists constant $C^{\prime}$ such that
(111) $\mathbb{P}\left\{r_{t}\leq C_{1}M\frac{6}{t-3}+(C_{1}\sqrt{\beta}+C_{2})C^{\prime}\left(\frac{3}{t-3}\right)^{\frac{\min\{\nu,1\}}{d}}\log^{\eta}\left(\frac{t}{3}\right)\right\}\geq 1-\delta.$
Hence, the convergence rate of GP-EI is $\mathcal{O}(t^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(t))$ in the noiseless case.
@@
By Lemma 7 in [17], there exists a constant $C^{\prime}>0$ such that for $\forall k\in\mathbb{N}$ and ${\bm{x}}_{i},i=1,\dots,t$, $\sigma_{i}({\bm{x}}_{i+1})\geq C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)$ holds at most $k$ times, where $\eta=\alpha$ if $\nu\leq 1$ and $\eta=0$ if $\nu>1$. Both $\nu>0$ and $\alpha\geq 0$ are parameters defining the properties of the kernels in assumptions 1-4 in [17].
Following the proof of Theorem 4.7, we let $k=\left[\frac{t}{3}\right]$. Then, there exists $k\leq t_{k}\leq 3k$ so that $y_{t_{k}}^{+}-y_{t_{k}+1}<\frac{2M}{k}$ and $\sigma_{t_{k}}({\bm{x}}_{t_{k}+1})<C^{\prime}k^{-\frac{\min\{\nu,1\}}{d}}\log^{\eta}(k)$. Applying Theorem 4.7, with probability $1-\delta$,
@@
Theorem 4.9 illustrates that the convergence rate in the noiseless case under the GP prior assumption is similar to that under the RKHS assumption in [17]. The analysis can be applied to Theorem 4.2 in the noiseless case as well, with a similar convergence rate. We note that both the SE and Matérn kernels satisfy assumptions (1)-(4) [17], with SE kernel obtained as $\nu\to\infty$.
The rate of decrease for the error bound with noise is given in the following theorem.
@@
Given $\delta\in(0,1)$, let $\beta=2\log(\frac{9c_{\alpha}}{\delta})$, $w=\sqrt{2\log(\frac{9}{2\delta})}$, and $c_{t}^{\sigma}=2\log(\frac{\pi^{2}t^{2}}{2\delta})$, where $c_{\alpha}=\frac{1+2\pi}{2\pi}$. Let $C_{1}=\frac{1}{\Phi(-w)}$ and $C_{2}=\frac{\phi(0)}{\Phi(-w)}+\sqrt{\beta}$. Then, for $t\geq\frac{4\log(\frac{3}{\delta})}{\log(2)}+4$, the error bound of GP-EI reduces at the rate
@@
Next, we prove that Theorem 4.7 provides tighter bounds than Theorem 4.2.
@@
We prove the noisy case, as the noiseless case is similar. To distinguish the parameters, we denote the $\beta$ in Theorem 4.7 as $\beta_{4.6}$ and $\beta$ in Theorem 4.2 as $\beta_{4.2}$. Hence, given the same probability $\delta$, $\beta_{4.6}=2\log(\frac{9c_{\alpha}}{\delta})$ and the $\beta_{4.2}=2\log(\frac{6}{\delta})$. To better compare the two theorems, we can rewrite both  (52) and (62) as
@@
(116) $C_{4}^{4.2}=\frac{\tau(\sqrt{\beta_{4.2}})}{\tau(-\sqrt{\beta_{4.2}})},\ C_{5}^{4.2}=\frac{\tau(\sqrt{\beta_{4.2}})}{\tau(-\sqrt{\beta_{4.2}})}(\sqrt{\beta_{4.2}}+\phi(0)),$
where the superscript indicates the corresponding theorems. Let $w_{4.6}=2\log(\frac{9}{2\delta})$. From (62), we have that
(117) $C_{4}^{4.6}=\frac{1}{\Phi(-\sqrt{w_{4.6}})},\ C_{5}^{4.6}=\frac{\sqrt{\beta_{4.6}}}{\Phi(-\sqrt{w_{4.6}})}+\frac{\phi(0)}{\Phi(-\sqrt{w_{4.6}})}+\sqrt{\beta_{4.6}}.$
Using $\beta_{4.2}>w_{4.6}$, we have that
(118) $C_{4}^{4.2}=\frac{\tau(\sqrt{\beta_{4.2}})}{\tau(-\sqrt{\beta_{4.2}})}>\frac{\tau(\sqrt{\beta_{4.2}})}{\Phi(-\sqrt{\beta_{4.2}})}>\frac{\sqrt{\beta_{4.2}}}{\Phi(-\sqrt{w_{4.6}})}>C_{4}^{4.6},$
where the first inequality uses Lemma 3.4, the second one uses Lemma 3.6, and the last one uses $\beta_{4.2}>1$. Notice that $w_{4.6}>2\log(\frac{9}{2})>3$. Then, using $\delta\in(0,1)$, $w_{4.6}<\beta_{4.2}$, and inequalities from (118), we have
(119) $C_{5}^{4.6}=$ $\frac{1}{\Phi(-\sqrt{w_{4.6}})}(\sqrt{\beta_{4.6}}+\phi(0)+\sqrt{\beta_{4.6}}\Phi(-\sqrt{w_{4.6}}))<\frac{1.042\sqrt{\beta_{4.6}}+\phi(0)}{\Phi(-\sqrt{\beta_{4.2}})}.$
It is easy to verify that $1.042\sqrt{\beta_{4.6}}+\phi(0)<\beta_{4.2}+\phi(0)\sqrt{\beta_{4.2}}$ by taking the square on both sides. Hence, by Lemma 3.4 and 3.6, we have
(120) $C_{5}^{4.6}<$ $\frac{1}{\Phi(-\sqrt{\beta_{4.2}})}(\beta_{4.2}+\phi(0)\sqrt{\beta_{4.2}})<\frac{\tau(\sqrt{\beta_{4.2}})(\sqrt{\beta_{4.2}}+\phi(0))}{\tau(-\sqrt{\beta_{4.2}})}=C_{5}^{4.2}.$
@@
An example of Proposition 4.14 is $\delta=0.1$. Then, $\beta_{4.2}=8.19$, $C_{4}^{4.2}=4632$ and $C_{5}^{4.2}=15103$. Meanwhile, $\beta_{4.6}=9.17$, $C_{1}=345$, and $C_{2}=141$. Thus, $C_{4}^{4.6}=345$, and $C_{5}^{4.6}=1187$. Additional values of $\delta$ are illustrated in Figure 5.
4.3 Improved error bound for RKHS objectives
In this section, we demonstrate how the analysis and techniques from Section 4.1 and 4.2 can be applied to RKHS objectives to improve existing best-known error bounds and convergence rates from [17] in the noiseless case. Formally, Assumptions 2.62.7 are valid in this section. The following lemma can be found in [30].